# Flow to Learn: Flow Matching on Neural Network Parameters

### v1

@@ 1 Introduction
Multiple approaches have been tried to generate weights capable of few-shot learning (FSL), motivated by its speed compared to conventional training. For instance, various diffusion-based approaches (Soro et al., 2024; Zhang et al., 2024; Wang et al., 2024) have been used to generate neural network weights. However, flexibility is limited by its restriction to Gaussian processes and a sluggish inference speed. More broadly, we may categorize this form of learning as meta-learning (Fifty et al., 2024; Hu et al., 2022; Zhmoginov et al., 2022), which aims to learn concepts from a few demonstrations. It is therefore natural that we have two evaluation settings: in-distribution tasks and out-of-distribution (OOD) tasks. With enough training and capacity, it’s clear meta-models (i.e. models trained on multiple data distributions) should excel at in-distribution tasks. However, generalization to novel tasks often presents a challenge to meta-learning and weight generation frameworks (Wang et al., 2024; Schürholt et al., 2024). Recently, Soro et al. (2024) goes to some extent to cross this generalization gap, however, we show that some improvements can be made to the approach. See Appendix A.1 for further related works.
@@ 2.2 Flow to Learn
Due to the intractable size of weight space, it is necessary for modeling to take place in latent space. We justify this design by appealing to work on the Lottery Ticket Hypothesis (Frankle & Carbin, 2019; Liu et al., 2024a) as well as the body of work on pruning (Cheng et al., 2024), which suggests that, like natural data, neural networks live on a low-dimensional manifold within its ambient space. We have two encoder variants, first is a variational autoencoder (VAE) (Kingma & Welling, 2022) set up as in Soro et al. (2024), and the second is the graph-based encoder (GE) of Kofinas et al. (2024) which takes into account permutation invariance of neural networks. As the latter models connections between layers, we only use the VAE for experiments involving subsets of weights.
@@
To condition our flow meta-model, we use a pre-trained CAML. Our choice is due to the extensive training of the CAML architecture on several datasets as well as the principled label encoding used in their approach. As we are only interested in encoding the support set, we consider judicious choices for the query set expected by CAML. For instance, in the Model Retrieval experiment below, we simply choose one random image for each class in the support set. In the FSL experiment, the choice is clear: the query set of each FSL task. The conditioning vector is incorporated by concatenating to the latent vector.
@@ A.1 Related Works
Following up on the work of meta-learning context, few-shot learning is a natural application of such meta-learning algorithms. An early example is Ravi & Larochelle (2017) who designed a meta-learner based on the computations in an LSTM cell. At each training example in the support set, their meta learner uses the losses and the gradients of the losses of the base learner (in addition to other information from previous training examples) to produce base learner parameters for the next training example. The loss of the base learner on the test examples in the support set is backpropagated through the meta learner’s parameters. Moreover, we may leverage the advancements in generative modeling for weight generation. As we mentioned, Lee et al. (2023) used transformers for in-context reinforcement learning, but we also see the works of Zhmoginov et al. (2022); Hu et al. (2022); Kirsch et al. (2024); Fifty et al. (2024) use transformers and foundation models. More similar to our method is the body of work on using diffusion models for weight generation (Du et al., 2023; Zhang et al., 2024; Wang et al., 2024; Soro et al., 2024). These methods vary in their approach, some leveraging a relationship between the gradient descent algorithm and the denoising step in diffusion models to design their meta-learning algorithm. Others rely on the modeling capabilities of conditioned latent diffusion models to learn the target distribution of weights. Most evaluations conducted were in-distribution tasks, i.e. tasks sampled from the same data distribution as the training tasks, hence, there is room to explore ways of adapting this approach for out-of-distribution tasks.

### v2

@@ 1 Introduction
Multiple approaches have been tried to generate weights capable of few-shot learning (FSL), motivated by its speed compared to conventional training. For instance, various diffusion-based approaches (Soro et al., 2024; Zhang et al., 2024; Wang et al., 2024) have been used to generate neural network weights. However, flexibility is limited by its restriction to Gaussian processes and a sluggish inference speed. More broadly, we may categorize this form of learning as meta-learning (Fifty et al., 2024; Hu et al., 2022b; Zhmoginov et al., 2022), which aims to learn concepts from a few demonstrations. It is therefore natural that we have two evaluation settings: in-distribution tasks and out-of-distribution (OOD) tasks. With enough training and capacity, it’s clear meta-models (i.e. models trained on multiple data distributions) should excel at in-distribution tasks. However, generalization to novel tasks often presents a challenge to meta-learning and weight generation frameworks (Wang et al., 2024; Schürholt et al., 2024). Recently, Soro et al. (2024) goes to some extent to cross this generalization gap, however, we show that some improvements can be made to the approach. See Appendix A.1 for further related works.
@@ 2.2 Flow to Learn
Due to the intractable size of weight space, it is necessary for modeling to take place in latent space. We justify this design by appealing to work on the Lottery Ticket Hypothesis (Frankle & Carbin, 2019; Liu et al., 2024a) as well as the body of work on pruning (Cheng et al., 2024), which suggests that, like natural data, neural networks live on a low-dimensional manifold within its ambient space. We use three encoder variants in this paper, first is a variational autoencoder (VAE) (Kingma & Welling, 2022) set up as in Soro et al. (2024), and the second is the graph-based encoder (GE) of Kofinas et al. (2024) which takes into account permutation invariance of neural networks, and third is based on the Learning on LoRAs (LoL) architecture (Putterman et al., 2024). Usage of each variant is dependent on experimental settings.
@@
To condition our flow meta-model, we use a pre-trained CAML (Fifty et al., 2024) adapated by Low-rank adaptation modules (LoRAs) (Hu et al., 2022a). Our choice is due to the extensive training of the CAML architecture on several datasets as well as the principled label encoding used in their approach. As we are only interested in an overall encoding of support images, we append a combining layer to aggregate the image latents. The conditioning vector is incorporated by concatenating to the latent vector.
@@ A.1 Related Works
Following up on the work of meta-learning context, few-shot learning is a natural application of such meta-learning algorithms. An early example is Ravi & Larochelle (2017) who designed a meta-learner based on the computations in an LSTM cell. At each training example in the support set, their meta learner uses the losses and the gradients of the losses of the base learner (in addition to other information from previous training examples) to produce base learner parameters for the next training example. The loss of the base learner on the test examples in the support set is backpropagated through the meta learner’s parameters. Moreover, we may leverage the advancements in generative modeling for weight generation. As we mentioned, Lee et al. (2023) used transformers for in-context reinforcement learning, but we also see the works of Zhmoginov et al. (2022); Hu et al. (2022b); Kirsch et al. (2024); Fifty et al. (2024) use transformers and foundation models. More similar to our method is the body of work on using diffusion models for weight generation (Du et al., 2023; Zhang et al., 2024; Wang et al., 2024; Soro et al., 2024). These methods vary in their approach, some leveraging a relationship between the gradient descent algorithm and the denoising step in diffusion models to design their meta-learning algorithm. Others rely on the modeling capabilities of conditioned latent diffusion models to learn the target distribution of weights. Most evaluations conducted were in-distribution tasks, i.e. tasks sampled from the same data distribution as the training tasks, hence, there is room to explore ways of adapting this approach for out-of-distribution tasks.